CdF2(s)⇄Cd2+(aq)+2F−(aq)

A saturated aqueous solution of CdF2 is prepared. The equilibrium in the solution is represented above. In the solution, [Cd2+]eq=0.0585M and [F−]eq=0.117M. Some 0.90MNaF is added to the saturated solution. Which of the following identifies the molar solubility of CdF2 in pure water and explains the effect that the addition of NaF has on this solubility? a. The molar solubility of CdF2 in pure water is 0.0585M, and adding NaF decreases this solubility because the equilibrium shifts to favor the precipitation of some CdF2.

b. The molar solubility of CdF2 in pure water is 0.0585M, and adding NaF has no effect on the solubility because only changes in temperature can increase or decrease the molar solubility of an ionic solid.

c.The molar solubility of CdF2 in pure water is 0.117M, and adding NaF decreases this solubility because the equilibrium shifts to favor the precipitation of some CdF2.

d. The molar solubility of CdF2 in pure water is 0.176M, and adding NaF increases this solubility because the Na+ ions displace the Cd2+ ions, causing the equilibrium to shift to favor the products.
Answer #1

Ksp = (0.0585)(0.117)2

=8.0×10-4  = 4s3

s = 0.0585 M

The molar solubility of CdF2 in pure water is 0.0585M, and adding NaF decreases this solubility because the equilibrium shifts to favor the precipitation of some CdF2.

(a) is correct option